# Supervised Manifold Learning for Functional Data

### v1

@@ 3.2 Geodesic Distance Estimation
To estimate the geodesic distance, we use the functional parallel transport unfolding (FPTU) developed in Tan et al., (2024), which is shown to be more accurate and robust than the traditional FIsomap. The first step is to construct an $n\times n$ weighting matrix $G$, where $G(i,j)=\|\widehat{X}_{i}-\widehat{X}_{j}\|_{L^{2}}$ if $\widehat{X}_{i}$ and $\widehat{X}_{j}$ are adjacent and $G(i,j)=\infty$ otherwise. To ensure that the graph induced by $G$ is connected so that the final manifold learning outcomes can be embedded into the same coordinate system, we utilize the minimum spanning tree idea in Dimeglio et al., (2014) to define the adjacency of $\{\widehat{X}_{i}\}_{i=1}^{n}$; see Section B of the Supplementary Material for details. For individuals $i$ and $j$, we define the geodesic path as the shortest path on $G$ between them, which is obtained by applying Dijkstra’s algorithm on $G$. The FIsomap uses the length of the geodesic path to estimate the geodesic distance, while the FPTU estimates the geodesic distance by the length of the unfolded geodesic path, which is obtained by the discrete parallel transport as follows.
We need to estimate the tangent space at each data point to perform the discrete parallel transport. The tangent space at a given point on the manifold is the best linear approximation of the manifold locally around that point. As PCA yields a linear projection of the data that is optimal in the sense of mean squared error, a natural idea of estimating tangent spaces is to utilize local PCA (Singer and Wu,, 2012). This argument also holds in the case of functional data. Therefore, to estimate $T_{i}\mathcal{M}\equiv T_{X_{i}}\mathcal{M}$, the tangent space at $X_{i}$, we consider the neighborhood $\mathcal{N}_{i,\rm{PCA}}$ that contains $k_{\mathrm{PCA}}$-nearest neighbors ($k_{\mathrm{PCA}}$-NN) of $\widehat{X}_{i}$ (note that $X_{i}$ is not available). We define the local empirical covariance function around $\widehat{X}_{i}$ as
@@
where $\widehat{\mu}_{i}(t)=\sum_{j\in\mathcal{N}_{i,\rm{PCA}}}\widehat{X}_{j}/k_{\rm{PCA}}$ is the empirical local mean around $\widehat{X}_{i}$. The estimator of $T_{i}\mathcal{M}$ is then defined as $\widehat{T}_{i}\mathcal{M}=\mathrm{span}\{\widehat{\phi}_{i1},\ldots,\widehat{\phi}_{id}\}$, where $\widehat{\phi}_{i1},\ldots,\widehat{\phi}_{id}$ are the first $d$ eigenfunctions of $\widehat{\Gamma}_{i}$. If the intrinsic dimension $d$ is unknown, we adapt the minimal neighborhood method (Facco et al.,, 2017) to estimate it; see Section B of the Supplementary Material for details.
[…]
@@ 3.3 Interpolation and Classification
Here, $\widehat{m}_{x}(t)=\sum_{j\in\mathcal{N}_{x,\rm{PCA}}}\widehat{X}_{j}/k_{\rm{PCA}}$ and $\mathcal{N}_{x,\rm{PCA}}$ contains $k_{\mathrm{PCA}}$-NN of $x$. We project $\widehat{X}_{i}$ on $\widehat{T}_{x}\mathcal{M}$, whose coordinate in terms of $\{\widehat{\phi}_{x1},\ldots,\widehat{\phi}_{xd}\}$ is $\widehat{\chi}_{i}=(\langle\widehat{X}_{i}-x,\widehat{\phi}_{x1}\rangle,\ldots,\langle\widehat{X}_{i}-x,\widehat{\phi}_{xd}\rangle)^{\top}$, for $i=1,\ldots,n$. The estimator of $\mu(x)$ is defined as
@@
where $\mathbb{W}=\mathrm{diag}\big{(}K_{h_{\mathrm{reg}}}(\|\widehat{X}_{1}-x\|_{L^{2}}),\ldots,K_{h_{\mathrm{reg}}}(\|\widehat{X}_{n}-x\|_{L^{2}})\big{)}$ with $K_{h}=K(\cdot/h)/h^{d}$, ^$\widehat{\mathbb{Z}}=(\widehat{Z}_{1},\ldots,\widehat{Z}_{n})^{\top}$, $e_{1}=(1,0,\ldots,0)^{\top}$ is a $d+1$-dimensional vector, and
@@
Finally, a classical multivariate classifier $\widehat{H}_{d}:\mathbb{R}^{d}\to\mathcal{Y}$, say the $k$-NN classifier, is trained on $\{\widehat{Z}_{i},Y_{i}\}_{i=1}^{n}$, and our functional classifier is defined as $\widehat{H}(\cdot)=\widehat{H}_{d}\{\widehat{\mu}(\cdot)\}$.
@@ 3.4 Tuning Parameters Selection
Our procedure contains several tuning parameters to be selected. In the first step of recovering functional data, we suggest using the plug-in bandwidth (Ruppert et al.,, 1995) for the ridged local linear estimator. Regarding geodesic distance estimation, following the suggestion of Tan et al., (2024), one may experiment several values around $n^{2/(d+2)}$ for $k_{\rm{PCA}}$. If the purpose is dimension reduction solely, one may test a few values for $\xi$ and apply the MDS on $\mathbb{D}_{\xi}$ to obtain the low-dimensional representations.
If the purpose is classification, it remains to choose $\xi$ in computing $\mathbb{D}_{\xi}$, which largely affects the manifold learning outcomes, and the functional regression bandwidth $h_{\rm{reg}}$, which determines the ability of the out-of-sample interpolation. These two parameters affect each other and are both crucial to classification performance. Therefore, a careful selection for these two parameters is needed. We propose a nested $L$-fold cross-validation (CV) procedure to select them. Specifically, we first estimate pairwise geodesic distances using the whole dataset. We then randomly split the dataset into $L$ parts denoted by $\mathcal{F}_{1},\ldots,\mathcal{F}_{L}$. Let $\mathcal{F}_{-\ell}$ denote the remaining dataset with $\mathcal{F}_{\ell}$ excluded. For a given $\mathcal{F}_{-\ell}$, we further randomly split it into $L$ parts, $\mathcal{S}_{1}^{\ell},\ldots,\mathcal{S}_{L}^{\ell}$, and denote by $\mathcal{S}_{-m}^{\ell}$ the remaining dataset with $\mathcal{S}_{m}^{\ell}$ excluded. The outer split samples $\mathcal{F}_{\ell}$ are used to select $\xi$, while the inner split samples $\mathcal{S}^{\ell}_{m}$ are used to select $h_{\rm{reg}}$. Given a classifier $\widehat{H}_{d}$ on $\mathbb{R}^{d}$, we define the CV loss for $\xi$ as
@@ 4 Theoretical Properties
Recall that we assumed that the manifold $\mathcal{M}$ is characterized by a single coordinate map $\mu:\mathcal{M}\to\mathcal{U}$ with an open $\mathcal{U}\in\mathbb{R}^{d}$. The coordinate map is not unique: For any diffeomorphism $\psi$, $\psi\{\mu(\cdot)\}$ is also a coordinate map. Since our manifold learning procedure aims to preserve the geodesic distance (up to a penalty for different labels), we are interested in coordinate maps that preserve the distance. Therefore, we assume that there exists a coordinate map $\mu$ that is isometric, i.e., $\|\mu(x_{1})-\mu(x_{2})\|_{\mathbb{R}^{d}}=d_{g}(x_{1},x_{2})$ for any $x_{1},x_{2}\in\mathcal{M}$. If such a $\mu$ exists, it is still not unique, because a rigid transformation applied to $\mu$ does not change the distance. In addition, identifying $\mu$ globally from the observed data alone without any assumptions is not possible, because the low-dimensional representations $Z_{i}=\mu(X_{i})$ are not observed. For $i=1,\ldots,n$, let $\widehat{\mu}(X_{i})=\widehat{Z}_{i}$ denote the manifold learning outcome. To avoid this dilemma, Chen and Müller, (2012) assumed beforehand that $E\{\sup_{i=1,\ldots,n}\|\widehat{\mu}(X_{i})-\mu(X_{i})\|_{\mathbb{R}^{d}}\}\to 0$ at a given rate, and investigated the convergence rate of their $\widehat{\mu}^{-1}$ at any given point. In contrast, we show that a more concrete convergence rate of $\sup_{i=1,\ldots,n}\|\widehat{\mu}(X_{i})-\mu(X_{i})\|_{\mathbb{R}^{d}}$ can be obtained under suitable conditions. All the proofs of the theoretical results can be found in the Supplementary Material.
@@
The assumption above requires that the distance between $\widehat{\mu}$ and $\mu$ at the sample points can be controlled by the difference of with-in-sample distances $\|\widehat{\mu}(X_{i})-\widehat{\mu}(X_{j})\|_{\mathbb{R}^{d}}$ and $\|\mu(X_{i})-\mu(X_{j})\|_{\mathbb{R}^{d}}$. If $\mu$ satisfies Assumption 1, $\psi\{\mu(\cdot)\}$ no longer satisfies this assumption with a rigid transformation $\psi$ on $\mathbb{R}^{d}$. Therefore, Assumption 1 can be seen as an idenfitiability condition for $\mu$. The idea of imposing such an assumption is that the convergence rate of the right hand side of (4.1) is tractable.
@@
Under Assumption 1 and Conditions (A1) to (A4), (B1) to (B5) in Section C.1 of the Supplementary Material, assuming that $\xi,\epsilon_{\rm{MDS}}\to 0$, and the number of vertices of any geodesic path on $G$ is uniformly bounded by $m$, we have
@@
where $h_{\rm{PCA}},h_{g},h_{d}$ and $\kappa_{s}$ are given in Section C.1 of the Supplementary Material.
@@
Under the same conditions as in Proposition 1 and Conditions (C1) and (C2) in Section C.2 of the Supplementary Material, suppose $h_{\rm{reg}}>h_{\rm{PCA}}$, for any $\epsilon>0$ and any given point $x\in\mathcal{M}$, we have
@@
Compared to Proposition 1, the convergence rate of $\widehat{\mu}$ at a given point includes the additional terms $h_{\rm{reg}}^{2}+h_{\rm{reg}}^{-\alpha-\epsilon d}J^{-\beta}$, which are the convergence rate of the bias of the local linear estimator in the tangent space (Lin and Yao,, 2021). Unlike the regression setting where the outcome variable is observed with error, the outcomes here are obtained by the supervised manifold learning procedure, and thus there is no variance part.
@@
$\widehat{H}_{k\rm{NN}}(X_{0})=\left\{\begin{aligned} &0,~{}\textrm{if }~{}\frac{1}{k}\sum_{i=1}^{k}Y_{(i)}\leq\frac{1}{2}\,,\\ &1,~{}\textrm{otherwise\,.}\end{aligned}\right.$ (4.2)
@@
Under the same conditions as Proposition 2 and Conditions (D1) to (D3) in Section D of the Supplementary Material, for all $\delta>0$, there exists $N_{\delta}\in\mathbb{N}$, such that if $n\gtrsim\max\{N_{\delta},\delta^{-(d+2)/d}\}$, $k=\lfloor\delta^{-2/d}\rfloor$, then
$\mathcal{R}(\widehat{H}_{k\rm{NN}})-\mathcal{R}(H^{*})\leq C\delta^{(1+\gamma)/d}\,,$ (4.3)
@@
Theorem 1 shows that the risk between $\widehat{H}_{k\rm{NN}}$ and $H^{*}$ can be arbitrarily small, i.e., $\widehat{H}_{k\rm{NN}}$ is asymptotically optimal. The value $N_{\delta}$ is related to the convergence rate of $\widehat{\mu}$ as in Proposition 2. However, a more specific form for $N_{\delta}$ is not available, because Proposition 2 is a statement of the convergence in probability rather than the almost sure convergence; see the proof for more details. If $N_{\delta}>\delta^{-(d+2)/d}$, then $\delta^{-(d+2)/d}=o(n)$, which implies that $n^{-(1+\gamma)/(d+2)}/\delta^{(1+\gamma)/d}=o(1)$. That is, the convergence rate of the bound in (4.3) is slower than $n^{-(1+\gamma)/(d+2)}$ provided that $N_{\delta}>\delta^{-(d+2)/d}$. If $N_{\delta}<\delta^{-(d+2)/d}$ otherwise, we can in fact choose $n\asymp\delta^{-(d+2)/d}$, which implies that $\mathcal{R}(\widehat{H}_{k\rm{NN}})-\mathcal{R}(H^{*})\leq Cn^{-(1+\gamma)/(d+2)}$, the minimax optimal rate of the $k$-NN classifier as shown in Gadat et al., (2016).

### v2

@@ 3.2 Geodesic Distance Estimation
To estimate the geodesic distance, we employ the functional parallel transport unfolding (FPTU) developed in Tan et al., (2024), which is shown to be more accurate and robust than the traditional FIsomap. The first step is to construct an $n\times n$ weighting matrix $G$, where $G(i,j)=\|\widehat{X}_{i}-\widehat{X}_{j}\|_{L^{2}}$ if $\widehat{X}_{i}$ and $\widehat{X}_{j}$ are adjacent, and $G(i,j)=\infty$ otherwise. To ensure that the graph induced by $G$ is connected so that the final manifold learning outcomes can be embedded into the same coordinate system, we utilize the minimum spanning tree idea in Dimeglio et al., (2014) to define the adjacency of $\{\widehat{X}_{i}\}_{i=1}^{n}$; see Section B of the Appendix for details. For individuals $i$ and $j$, we define the geodesic path as the shortest path on $G$ between them, obtained by applying Dijkstra’s algorithm on $G$. The FIsomap uses the length of the geodesic path to estimate the geodesic distance, while the FPTU estimates the geodesic distance by the length of the unfolded geodesic path, which is obtained through the discrete parallel transport as follows.
We need to estimate the tangent space at each data point to perform the discrete parallel transport. The tangent space at a given point on the manifold serves as the best linear approximation of the manifold locally around that point. As PCA yields a linear projection of the data that is optimal in the sense of mean squared error, a natural idea of estimating tangent spaces is to utilize the local PCA (Singer and Wu,, 2012). This reasoning also applies in the context of functional data. Therefore, to estimate $T_{i}\mathcal{M}\equiv T_{X_{i}}\mathcal{M}$, the tangent space at $X_{i}$, we consider the neighborhood $\mathcal{N}_{i,\rm{PCA}}$ that contains $k_{\mathrm{PCA}}$-nearest neighbors ($k_{\mathrm{PCA}}$-NN) of $\widehat{X}_{i}$ (note that $X_{i}$ is not available). We define the local empirical covariance function around $\widehat{X}_{i}$ as
@@
where $\widehat{\mu}_{i}(t)=\sum_{j\in\mathcal{N}_{i,\rm{PCA}}}\widehat{X}_{j}/k_{\rm{PCA}}$ is the empirical local mean around $\widehat{X}_{i}$. The estimator of $T_{i}\mathcal{M}$ is then defined as $\widehat{T}_{i}\mathcal{M}=\mathrm{span}\{\widehat{\phi}_{i1},\ldots,\widehat{\phi}_{id}\}$, where $\widehat{\phi}_{i1},\ldots,\widehat{\phi}_{id}$ are the first $d$ eigenfunctions of $\widehat{\Gamma}_{i}$. If the intrinsic dimension $d$ is unknown, we adapt the minimal neighborhood method (Facco et al.,, 2017) to estimate it; see Section B of the Appendix for details.
Let $\widehat{\Phi}_{ij}$ denote a $d\times d$ matrix with the $(k,s)$th entry $\widehat{\Phi}_{ij}(k,s)=\langle\widehat{\phi}_{ik},\widehat{\phi}_{js}\rangle$, where $\langle\cdot,\cdot\rangle$ is the inner product in $L^{2}(\mathcal{T})$. It can be shown that $R_{j,i}=VU^{\top}$ provides an approximate parallel transport from $\widehat{T}_{i}\mathcal{M}$ to $\widehat{T}_{j}\mathcal{M}$, where $\widehat{\Phi}_{ij}=U\Sigma V^{\top}$ is the singular value decomposition of $\widehat{\Phi}_{ij}$ (Tan et al.,, 2024). This means that, for a vector $u=(u_{1},\ldots,u_{d})^{\top}\in\widehat{T}_{i}\mathcal{M}$ in terms of the basis $\{\widehat{\phi}_{i1},\ldots,\widehat{\phi}_{id}\}$, the vector $\widehat{R}_{j,i}(u_{1},\ldots,u_{d})^{\top}$ is the parallel transported $u$ in $\widehat{T}_{j}\mathcal{M}$ under the basis $\widehat{\Phi}_{j}$. For a geodesic path $(i=i_{0},i_{1},\ldots,i_{m-1},i_{m}=j)$ in $G$, we parallel transport all the edges on the path to the last tangent space $\widehat{T}_{j}\mathcal{M}$ and then aggregate them, which yields the unfolded geodesic. The Euclidean norm of the unfolded geodesic provides a robust approximation of the geodesic distance between $i$ and $j$. Specifically, we first project $\widehat{V}_{i_{k}}=\widehat{X}_{i_{k-1}}-\widehat{X}_{i_{k}}$ to the tangent space $\widehat{T}_{i_{k}}\mathcal{M}$, for $k=1,\ldots,m$. Let $v_{i_{k}}=(\langle\widehat{V}_{i_{k}},\widehat{\phi}_{i_{k}1}\rangle,\ldots,\langle\widehat{V}_{i_{k}},\widehat{\phi}_{i_{k}d}\rangle)^{\top}$ be the coordinates of the projected $\widehat{V}_{i_{k}}$. It follows that $v_{i_{k},m}=\widehat{R}_{i_{m},i_{m-1}}\cdots\widehat{R}_{i_{k+1},i_{k}}v_{i_{k}}$ is the parallel transported $v_{i_{k}}$ in $\widehat{T}_{j}\mathcal{M}$, for $k=1,\ldots,m-1$. For $k=m$, we define $v_{i_{k},m}=v_{j,m}=v_{j}$, which is already in $\widehat{T}_{j}\mathcal{M}$ and no parallel transport is needed. Finally, the aggregated vector $v_{i,m}=\sum_{k=1}^{m}v_{i_{k},m}$ in $\widehat{T}_{j}\mathcal{M}$ can be seen as the unfolded geodesic from $\widehat{X}_{i}$ to $\widehat{X}_{j}$. As the Euclidean norm of the unfolded geodesic from $\widehat{X}_{j}$ to $\widehat{X}_{i}$ is generally different from that of $v_{i,m}$, we compute the average of these two as $\widehat{d}_{\mathcal{M}}(X_{i},X_{j})$, our estimator of $d_{\mathcal{M}}(X_{i},X_{j})$.
@@ 3.3 Interpolation and Classification
Here, $\widehat{m}_{x}(t)=\sum_{j\in\mathcal{N}_{x,\rm{PCA}}}\widehat{X}_{j}/k_{\rm{PCA}}$ and $\mathcal{N}_{x,\rm{PCA}}$ contains $k_{\mathrm{PCA}}$-NN of $x$. We project $\widehat{X}_{i}$ onto $\widehat{T}_{x}\mathcal{M}$, whose coordinates in terms of $\{\widehat{\phi}_{x1},\ldots,\widehat{\phi}_{xd}\}$ are $\widehat{\chi}_{i}=(\langle\widehat{X}_{i}-x,\widehat{\phi}_{x1}\rangle,\ldots,\langle\widehat{X}_{i}-x,\widehat{\phi}_{xd}\rangle)^{\top}$, for $i=1,\ldots,n$. The estimator of $\mu(x)$ is defined as
@@
where $\mathbb{W}=\mathrm{diag}\big(K_{h_{\mathrm{reg}}}(\|\widehat{X}_{1}-x\|_{L^{2}}),\ldots,K_{h_{\mathrm{reg}}}(\|\widehat{X}_{n}-x\|_{L^{2}})\big)$ with $K_{h}=K(\cdot/h)/h^{d}$, $\widehat{\mathbb{Z}}=(\widehat{Z}_{1},\ldots,\widehat{Z}_{n})^{\top}$, $e_{1}=(1,0,\ldots,0)^{\top}$ is a $d+1$-dimensional vector, and
@@
[…]
@@ 3.4 Tuning Parameters Selection
Our procedure involves selecting several tuning parameters. In the first step of recovering functional data, we suggest using the plug-in bandwidth (Ruppert et al.,, 1995) for the ridged local linear estimator. Regarding geodesic distance estimation, following the suggestion of Tan et al., (2024), one may experiment several values around $n^{2/(d+2)}$ for $k_{\rm{PCA}}$. If the purpose is dimension reduction solely, one may test a few values for $\xi$ and apply the MDS on $\mathbb{D}_{\xi}$ to obtain the low-dimensional representations.
If the purpose is classification, it remains to choose $\xi$ in computing $\mathbb{D}_{\xi}$, which significantly affects the manifold learning outcomes, and the functional regression bandwidth $h_{\rm{reg}}$, which determines the ability of the out-of-sample interpolation. These two parameters affect each other and are both crucial to classification performance. Therefore, a careful selection for these two parameters is needed. We propose a nested $L$-fold cross-validation (CV) procedure to select them. Specifically, we first estimate pairwise geodesic distances using the whole dataset. We then randomly split the dataset into $L$ parts denoted by $\mathcal{F}_{1},\ldots,\mathcal{F}_{L}$. Let $\mathcal{F}_{-\ell}$ denote the remaining dataset with $\mathcal{F}_{\ell}$ excluded. For a given $\mathcal{F}_{-\ell}$, we further randomly split it into $L$ parts, $\mathcal{S}_{1}^{\ell},\ldots,\mathcal{S}_{L}^{\ell}$, and denote by $\mathcal{S}_{-m}^{\ell}$ the remaining dataset with $\mathcal{S}_{m}^{\ell}$ excluded. The outer split samples $\mathcal{F}_{\ell}$ are used to select $\xi$, while the inner split samples $\mathcal{S}^{\ell}_{m}$ are used to select $h_{\rm{reg}}$. Given a classifier $\widehat{H}_{d}$ on $\mathbb{R}^{d}$, we define the CV loss for $\xi$ as
@@ 4 Theoretical Properties
Recall that we assume that the manifold $\mathcal{M}$ is characterized by a single coordinate map $\mu:\mathcal{M}\to\mathcal{U}$ with an open $\mathcal{U}\in\mathbb{R}^{d}$. The coordinate map is not unique: For any diffeomorphism $\psi$, $\psi\{\mu(\cdot)\}$ is also a coordinate map. Since our manifold learning procedure aims to preserve the geodesic distance (up to a penalty for different labels), we are interested in coordinate maps that preserve the distance. Therefore, we assume that there exists a coordinate map $\mu$ that is isometric, i.e., $\|\mu(x_{1})-\mu(x_{2})\|_{\mathbb{R}^{d}}=d_{g}(x_{1},x_{2})$ for any $x_{1},x_{2}\in\mathcal{M}$. If such a $\mu$ exists, it is still not unique, because a rigid transformation applied to $\mu$ does not change the Euclidean distance. In addition, identifying $\mu$ globally from the observed data alone without any assumptions is not possible, because the low-dimensional representations $Z_{i}=\mu(X_{i})$ are not observed. For $i=1,\ldots,n$, let $\widehat{\mu}(X_{i})=\widehat{Z}_{i}$ denote the manifold learning outcome. To avoid this dilemma, Chen and Müller, (2012) assumed beforehand that $E\{\sup_{i=1,\ldots,n}\|\widehat{\mu}(X_{i})-\mu(X_{i})\|_{\mathbb{R}^{d}}\}\to 0$ at a given rate, and investigated the convergence rate of their $\widehat{\mu}^{-1}$ at any given point. In contrast, we show that a more concrete convergence rate of $\sup_{i=1,\ldots,n}\|\widehat{\mu}(X_{i})-\mu(X_{i})\|_{\mathbb{R}^{d}}$ can be obtained under suitable conditions. All the proofs of the theoretical results can be found in Sections C and D of the Appendix.
@@
The assumption above requires that the distance between $\widehat{\mu}$ and $\mu$ at the sample points can be controlled by the difference of with-in-sample distances $\|\widehat{\mu}(X_{i})-\widehat{\mu}(X_{j})\|_{\mathbb{R}^{d}}$ and $\|\mu(X_{i})-\mu(X_{j})\|_{\mathbb{R}^{d}}$. If $\mu$ satisfies Assumption 1, $\psi\{\mu(\cdot)\}$ no longer satisfies this assumption with a rigid transformation $\psi$ on $\mathbb{R}^{d}$. Therefore, Assumption 1 can be seen as an identifiability condition for $\mu$. The idea of imposing such an assumption is that the convergence rate of the right hand side of (4.1) is tractable. It is not necessary if the convergence rate of the left hand side of (4.1) is given.
@@
Under Assumption 1 and Conditions (A1) to (A4), (B1) to (B5) in Section C.1 of the Appendix, assuming that $\xi,\epsilon_{\rm{MDS}}\to 0$, and the number of vertices of any geodesic path on $G$ is uniformly bounded by $m$, we have
@@
where $h_{\rm{PCA}},h_{g},h_{d}$ and $\kappa_{s}$ are given in Section C.1 of the Appendix.
@@
Under the same conditions as in Proposition 1 and Conditions (C1) and (C2) in Section C.2 of the Appendix, suppose $h_{\rm{reg}}>h_{\rm{PCA}}$, for any $\epsilon>0$ and any given point $x\in\mathcal{M}$, we have
@@
Compared to Proposition 1, the convergence rate of $\widehat{\mu}$ at a given point includes the additional terms $h_{\rm{reg}}^{2}+h_{\rm{reg}}^{-\alpha-\epsilon d}J^{-\beta}$, which are the convergence rate of the bias of the local linear estimator in the tangent space (Lin and Yao,, 2021). Unlike the regression setting where the outcome variable is observed with error, the outcomes here are obtained by the supervised manifold learning procedure, and thus there is no variance component.
@@
[…]
@@
Under the same conditions as Proposition 2 and Conditions (D1) to (D3) in Section D of the Appendix, for all $\delta>0$, there exists $N_{\delta}\in\mathbb{N}$, such that if $n\gtrsim\max\{N_{\delta},\delta^{-(d+2)/d}\}$, $k=\lfloor\delta^{-2/d}\rfloor$, then
$\mathcal{R}(\widehat{H}_{k-\rm{NN}})-\mathcal{R}(H^{*})\leq C\delta^{(1+\gamma)/d}\,$, (4.3)
@@
Theorem 1 shows that the risk between $\widehat{H}_{k-\rm{NN}}$ and $H^{*}$ can be arbitrarily small, i.e., $\widehat{H}_{k-\rm{NN}}$ is asymptotically optimal. The value $N_{\delta}$ is related to the convergence rate of $\widehat{\mu}$ as in Proposition 2. However, a more specific form for $N_{\delta}$ is not available, because Proposition 2 is a statement of the convergence in probability rather than the almost sure convergence; see the proof for more details. If $N_{\delta}>\delta^{-(d+2)/d}$, then $\delta^{-(d+2)/d}=o(n)$, which implies that $n^{-(1+\gamma)/(d+2)}/\delta^{(1+\gamma)/d}=o(1)$. That is, the convergence rate of the bound in (4.3) is slower than $n^{-(1+\gamma)/(d+2)}$ provided that $N_{\delta}>\delta^{-(d+2)/d}$. If $N_{\delta}<\delta^{-(d+2)/d}$ otherwise, we can in fact choose $n\asymp\delta^{-(d+2)/d}$, which implies that $\mathcal{R}(\widehat{H}_{k-\rm{NN}})-\mathcal{R}(H^{*})\leq Cn^{-(1+\gamma)/(d+2)}$, the minimax optimal rate of the $k$-NN classifier as shown in Gadat et al., (2016).